Force Application (Objectives): Library and Analysis

John Kingston AIAI University of Edinburgh

Contents

1 Introduction 4

2 Goals 4

3 Centers of Gravity 6

4 Constraints 9

5 Summary 11

A CinC’s objectives 11 A.1 Leadership . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 A.2 Key Production . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 A.3 National Infrastructure . . . . . . . . . . . . . . . . . . . . . . . . 12 A.4 National Population . . . . . . . . . . . . . . . . . . . . . . . . . 13 A.5 Fielded Forces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

B Air Objectives 14 B.1 Leadership . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 B.2 Key Production . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 B.3 National Infrastructure . . . . . . . . . . . . . . . . . . . . . . . . 15 B.4 Fielded Forces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

C Task Objectives 16

C.1 Leadership . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 C.1.1 YOC forces . . . . . . . . . . . . . . . . . . . . . . . . . . 16 C.1.2 Blue forces . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 C.2 Key Production . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 C.3 National Infrastructure . . . . . . . . . . . . . . . . . . . . . . . . 18 C.4 National Population . . . . . . . . . . . . . . . . . . . . . . . . . 18 C.5 Fielded Forces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

C.5.1 YOC forces . . . . . . . . . . . . . . . . . . . . . . . . . . 18 C.5.2 YOC supply, storage & repair facilities . . . . . . . . . . . 19 C.5.3 Blue forces . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

C.6 High level objectives . . . . . . . . . . . . . . . . . . . . . . . . . 20 C.7 Unclassified . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

D Goals 20 D.1 Goals identified . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 D.2 Ordering and definitions of goals . . . . . . . . . . . . . . . . . . . 23 D.3 Goals and suggested opposite goals . . . . . . . . . . . . . . . . . 23

E Centers of Gravity 26

F Constraints 39

List of Tables

1 Goals, ordered by effect . . . . . . . . . . . . . . . . . . . . . . . . 5 2 Goals and their opposites . . . . . . . . . . . . . . . . . . . . . . . 6 3 Constraints: CinC Objectives Level . . . . . . . . . . . . . . . . . 10 4 Goals, ordered by effect . . . . . . . . . . . . . . . . . . . . . . . . 24 5 Goals and their opposites . . . . . . . . . . . . . . . . . . . . . . . 25 6 Constraints: Air Objectives Level . . . . . . . . . . . . . . . . . . 41

List of Figures

1 Centers of Gravity related to Transportation . . . . . . . . . . . . 8 2 CoGS: Leadership . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3 CoGS: Key Production 1: General . . . . . . . . . . . . . . . . . . 28 4 CoGS: Key Production 2: Military . . . . . . . . . . . . . . . . . 29 5 CoGS: National Infrastructure 1: Transportation . . . . . . . . . 30 6 CoGS: National Infrastructure 2: Communications . . . . . . . . . 31 7 CoGS: National Infrastructure 3: Fuel & Power . . . . . . . . . . 32 8 CoGS: National Population . . . . . . . . . . . . . . . . . . . . . 33 9 CoGS: Fielded Forces 1: Special Weapons . . . . . . . . . . . . . 34 10 CoGS: Fielded Forces 2: Land forces . . . . . . . . . . . . . . . . 35 11 CoGS: Fielded Forces 3: Naval forces . . . . . . . . . . . . . . . . 36 12 CoGS: Fielded Forces 4: Air forces . . . . . . . . . . . . . . . . . 37 13 CoGS: Fielded Forces 5: Logistics . . . . . . . . . . . . . . . . . . 38

3

1 Introduction

This document analyses a list of CinC objectives, air objectives, and task objectives which have been used to support planning in recent military exercises. The list, which can be found in Appendices A to C, has been gathered and categorized by staff at Checkmate.

This document describes these objectives using the following template:

Achieve Goal on Center of Gravity with (optional) constraint.

The goals, centers of gravity, and constraints are listed separately. Some observations are then made.

20 JANUARY 1996: This is the first draft of this document. It therefore includes a number of assumptions, and probably some misunderstandings, which ought to be verified by the expert planners. Please do not publish this document in any form until verification has been carried out. Verification is especially needed in the key areas of this document: the table of goals and their effects in table D.1, and the hierarchies of centers of gravity shown in appendix E. In these hierarchies, a number of unlabelled boxes exist; these indicate that the hierarchy is not expected to be complete unless at least one center of gravity is defined at this level of the hierarchy.

N.B. Where the task objectives named actual locations, the names have been changed.

2 Goals

A number of goals were identified, at three different levels. A sample is shown below; the full list can be found in Appendix D.1.

• CinC Objectives

– National Infrastructure

∗ Deny use;

∗ Interrupt;

∗ Obtain;

∗ Permit use;

∗ Sustain

• Air Objectives

– National Infrastructure

∗ Drive (i.e. compel to move); ∗ Disrupt;

∗ Attack;

∗ Deploy.

• Task Objectives

– National Infrastructure

∗ Disrupt;

∗ Destroy.

Observations

1. The goals imply different levels of effect against centers of gravity. For example, deny use appears to be a stronger effect than disrupt. A table has been prepared which shows suggested orderings of goals, from the least effect at the top to the greatest effect at the bottom. A sample can be seen in Table 1; the full table can be found in table 4.

Goal  Definition

Leadership

Disrupt  Temporarily reduce effectiveness
Degrade  Permanently reduce effectiveness
Isolate Reduce effectiveness to zero until further notice
Destroy  Permanently reduce effectiveness to zero

Key production

Disrupt  Temporarily reduce production
Attrit  Permanently reduce production
Destroy  Permanently stop production

Infrastructure

Disrupt  Reduce the flow
Interrupt  Temporarily stop the flow
Deny use  Stop the flow until further notice

Table 1: Goals, ordered by effect

2. The planners identify Blue (friendly) centers of gravity as well as YOC centers of gravity. This can be seen from the use of goals such as permit and sustain as well as deny use and disrupt. It seems likely that, for every goal which can be performed against YOC centers of gravity, there is an equal and opposite goal to support Blue centers of gravity. Table 5 attempts to present a set of opposite goals for each goal identified in Table 4; that is, if the YOC forces tried to apply the goal named in the left hand column to a Blue centre of gravity, the goal in the right hand column

Goal  Opposite

Leadership

Disrupt Put into order
Degrade  Improve
Isolate  Connect
Destroy  Create

Key production

Disrupt Put into order
Destroy Build/Create

Infrastructure

Disrupt Put into order
Interrupt  Restart
Deny use Permit use

Table 2: Goals and their opposites

provide the opposite effect. A sample of entries from table 5 is shown in table 2.

This table provides the planners with alternatives to preventing disruption/destruction of Blue centers of gravity; it could also be used when Blue forces want to make use of centers of gravity (e.g. road networks) previously held by YOC forces.

3 Centers of Gravity

The centers of gravity which have been elicited from the sets of objectives have been categorized into hierarchies. These hierarchies are shown in appendix E; an example is shown in Figure 1. The aim of producing these hierarchies is to permit a near-exhaustive approach to identification of centers of gravity. By working from the top of a hierarchy downwards, planners can easily identify if they have considered all relevant categories of centers of gravity.

The 5 top level centers of gravity in these hierarchies are Leadership, Key

Production, National Infrastructure, National Population and Fielded

Forces. These correspond to Col. Warden’s ”5 rings” approach to identification of centers of gravity. The approach which has been taken to developing these hierarchies is top-down extrapolation; that is, if there are several centers of gravity at the same level of a hierarchy, and one of them is decomposed in a certain way, then the others are assumed to be decomposable in the same way. An example of this can be found in Figure 1, which shows centers of gravity connected with transportation. Since road, rail, inland waterways, coastal shipping and civil air transport are all transportation networks, then the hierarchy assumes that the key centers of gravity for all these types of transportation can be categorized as nodes, choke points or segments.

A key component of the hierarchies which contributes to the goal of identifying all relevant centers of gravity is the identification of inter-hierarchy links. These show where a center of gravity in a different hierarchy is relevant to a center of gravity in this hierarchy. The name of the related hierarchy is displayed in the bottom half of the CoG node. For example, Figure 6, which shows centers of gravity related to communication, includes ”Couriers” as a center of gravity. Since communication by courier is dependent on a good transportation network, an inter-hierarchy link to ”Transportation” is shown as the decomposition of the ”Couriers” center of gravity.

Observations

1. The centers of gravity which are identified in the different types of objectives correlate fairly closely with the different levels of the hierarchies. Most of the centers of gravity identified in the CinC objectives appear in the third level of the hierarchy (e.g. national roadways; most centers of gravity identified in the air objectives appear in the fourth level of the hierarchy (e.g choke points); and most centers of gravity identified in the task objectives appear in the fifth level of the hierarchy (e.g. bridges).

2. While the planners do identify Blue Centers of Gravity as well as YOC Centers of Gravity, there is evidence that Blue Centers of Gravity are not considered in the same level of detail. The centers of gravity identified for Blue in the CinC objectives (e.g. “sustain forces for a minimum of 45 days”) and in the air objectives (e.g. “deploy forces”) belong to the first or second levels of the hierarchies of centers of gravity. In addition, the goals applied to these centers of gravity imply a higher level of abstraction; for example, sustaining a force implies supplying all the needs of that force, which implicitly requires a complete infrastructure to be established for that force (bringing in fuel, taking out casualties, etc). Infrastructure is one of Col. Warden’s ”5 rings”, and can therefore be found at the top level of the hierarchy of centers of gravity. Similarly, deploying a force requires considerable attention to be paid to the transportation requirements of that force^{1} including but not limited to military air transport; transportation is a 2nd level center of gravity. It would seem sensible (if time is available) for the planners to analyse Blue centers of gravity to the same level of detail as YOC centres of gravity (if possible), to the level of identifying likely targets for any YOC planners who might be conducting a similar exercise.

1This assumes that deploying a force refers to the movement of a force into the battleground, as opposed to allocating new units to the theatre.

Figure 1: Centers of Gravity related to Transportation 8

4 Constraints

Many CinC objectives and Air objectives, and some task objectives, have associated constraints. These constraints may be:

• geographical (e.g. “Destroy SAMs beyond 100km of the DMZ”);

• temporal (e.g. “Sustain (Blue) forces for a minimum of 45 days”);

• categorical (e.g. “Deny use of national roadways for the transport of military equipment”);

• capability (e.g. “Deploy (Blue) forces capable of deterring YOC aggression”).

A list of most of the identified constraints was prepared^{2}. A selection of identified constraints is shown below. The full list can be seen in tables F and 6.

Observations

The constraints are diverse, but a few patterns can be seen:

• The distinction between military and civilian is frequently drawn. Typically, the goals specified will have stronger effects on the military than on the civilian population. For example, objectives exist to deny use of national roadways, railways and waterways for transporting military equipment, but only to disrupt the flow of goods. However, in some cases (e.g. disrupting communication networks), it is necessary to cause equal disruption to both the military and the civilian populations.

• Another distinction which appears to be important is the distinction between weapons of mass destruction and other weapons. This is reflected in the hierarchies of centers of gravity, where ”special weapons” (i.e. nuclear/ biological/ chemical weapons and ballistic missiles) are treated separately from other weapons.

• In the higher level objectives, the majority of geographical constraints are concerned with fielded forces, and their locations in relation to battlegrounds and demilitarized zones. However, there are also some concerns related to the whole region, typically connected with evacuation of U.S. civilians and other non-combatants.

2Those which were deemed to be linked to high level centers of gravity were eliminated.

Geographical

US citizens  in the region
US forces  in the region

 YOC aggression  against B/G

YOC forces  north of Desired Area

Temporal

YOC economic centers

to slow the flow of military resources i.e. until the flow of military resources is sufficiently slow

Categorical

YOC political control
YOC military control

YOC armed forces

YOC political control
YOC military control
   electricity
petroleum products
national television systems
national radio systems
national computer systems

YOC population

national roadways
national railways
national waterways

Transport of military equipment

national roadways
national railways
national waterways

Transport of goods

Table 3: Constraints: CinC Objectives Level

5 Summary

This document has analyzed CinC objectives, air objectives and task objectives, and has discerned a number of interesting patterns. These patterns have been reflected in tables of goals, and in hierarchies of centers of gravity. If these tables and hierarchies can be fully verified, they will provide useful domain knowledge for the expert task of identifying objectives, which is a key part of the planning process, and which is being modelled in detail in [the full document which AIAI are writing]. It is also possible that these tables and hierarchies may be of use to the planners in their current format.

A CinC’s objectives

The code numbers at the beginning of each objective show which higher level objectives are addressed by the lower level objectives; for example, air objective B5.1 fulfils CinC objective B5.

This document describes these objectives using the following template:

Achieve Goal on Center of Gravity with optional constraint.

Some objectives are rephrased to enable them to fit this template. The goal and the Center of Gravity are highlighted by using small capitals and teletype font respectively.

A.1 Leadership

A1: Isolate and neutralize YOC political control of armed forces

A2: Isolate and neutralize YOC political leadership from the population

A3: Ensure regional stability by reducing Yellowland Orangeland Cooperative (YOC). This is a national or political objective which should really be considered at a higher level of command than the other CinC objectives. However, it has been rephrased to:

     Reduce Yellowland Orangeland Cooperative (YOC) to ensure regional

stability

A4: Provide for Stability of the Region. This too is a national or political objective; it is too loosely defined for rephrasing to be possible.

A5: Protect the lives of US citizens and forces in the region. This is probably better categorized as a National Population objective.

A6: Isolate and neutralize YOC military control of armed forces A7: Isolate and neutralize YOC military’s effect on population. For consistency

with objective A6, this has been rephrased to:

Isolate and neutralize YOC military control on population

A.2 Key Production

B1: Separate YOC military structure from primary electrical sources. In order to match objective B2, this has been rephrased to: Deny use of electricity to YOC military structure.

B2: Deny use of primary electrical sources to YOC population. Rephrased^{3} to:

Deny use of electricity to YOC population B3: Neutralize YOC petroleum production supporting warfighting effort B4: Deny use of petroleum products to YOC population

B5: Neutralize production of chemical products

B6: Prevent chemical products from becoming weapons of mass destruction. This has been rephrased to: Deny use of chemical products for the manufacture of weapons of mass destruction.

B7: Neutralize YOC special weapons capability and delivery means

B8: Defend against Weapons of Mass Destruction. The current phrasing of this objective is identical to objective E7; it has therefore been removed from this section. Air objectives B8.1, B8.2 and B8.3 are renumbered to E7.1, E7.2 and E7.3.

A.3 National Infrastructure

C1: Deny use of national roadways for transport of military equipment C2: Deny use of national railways for transport of military equipment C3: Deny use of national waterways for transport of military equipment C4: Deny use of national communications nets to transport military infor-

3The term ”electrical production”, which appears in the Air Objectives (e.g. B2.1), is more general than ”primary electrical sources”. However, ”electricity” is a more general term still, which is at the same level as ”petroleum products”.

mation C5: Interrupt the flow of goods on national roadways C6: Interrupt the flow of goods on national railways C7: Interrupt the flow of goods on national waterways C8: Interrupt the flow of information on national communication networks C9: Build sufficient combat power for future offensive operations. This has been

rephrased to:

Obtain sufficent Blue Forces for future offensive operations. C10: Ensure Freedom of Navigation. This has been rephrased to:

Permit safe access to national waterways for Blue shipping. C11: Sustain (Blue) Forces for Minimum of 45 Days

A.4 National Population

D1: Deny use of national television systems to transmit military infor-

mation D2: Deny use of national radio systems to transmit military information D3: Deny use of national computer systems to transmit military informa-

tion D4: Deny use of national television systems to influence YOC population D5: Deny use of national radio systems to influence YOC population D6: Deny use of national computer systems to influence YOC population

D7: Conduct NEO of US personnel and other volunteer Foreign Nationals

D8: Neutralize YOC economic centers to slow the flow of military resources D9: Deny access to YOC economic centers. This has been rephrased to:

Deny use of national transportation networks leading to YOC economic centers. In this form, it can be seen that this is an infrastructure objective rather than a national population objective.

D10: Employ C2W methodologies.

Use C2W methodologies to influence YOC population.

This is a high level objective; in order to match the level of detail of the other objectives, individual C2W methods should be specified.

A.5 Fielded Forces

E1: Deter YOC aggression against B/G

E2: Deter Yellowland aggression

E3: Conduct Land, Sea and Air operations to restore the sovereignty of B/G. Like A3, this is a national or political objective which should really be considered at a higher level of command than the other CinC objectives. However, it has been rephrased to: Restore the sovereignty of B/G by fielding military forces.

E4: Stop YOC forces north of Desired Area E5: Defend B/G from YOC Attack E6: Defend Greyland from YOC Attack E7: Defend Against Weapons of Mass Destruction. Rephrased to:

Defend Blue forces Against Weapons of Mass Destruction

E8: Disrupt YOC Special Forces Capability

B Air Objectives

B.1 Leadership

A1.1: Disrupt YOC political control of armed forces A1.2: Destroy YOC political control of armed forces

A1.3: Disrupt and Degrade YOC intelligence capability
A1.4: Isolate YOC National Command Authority

A2.1: Disrupt YOC political control of population A2.5: Destroy YOC political control of population A6.1: Disrupt YOC military control of armed forces A6.2: Destroy YOC military control of armed forces

A6.3: Disrupt and Degrade YOC intelligence capability

A7.1: Disrupt YOC military control of population A7.2: Destroy YOC military control of population

B.2 Key Production

B1.1: Disrupt key electrical facilities providing power to YOC military forces B1.2: Destroy electrical facilities providing power to YOC mil-

itary forces

B2.1: Disrupt electrical production to YOC
B2.2: Destroy electrical production to YOC

B3.1: Disrupt petroleum production providing fuel to YOC military forces B3.2: Destroy petroleum production providing fuel to YOC military forces B4.1: Disrupt petroleum production throughout YOC B4.2: Destroy petroleum production throughout YOC B5.1: Disrupt chemical production facilities B5.2: Destroy chemical production facilities B7.1: Destroy strategic targets (pol-mil C2 and NBC/TBM capability)

B.3 National Infrastructure

C1.1: Drive YOC into desired chokepoints and restricted LOCs
C1.2: Disrupt movement of YOC forward corps and committed 2nd echelon
     forces
C2.1: Drive YOC into desired chokepoints and restricted LOCs
C2.2: Disrupt movement of YOC forward corps and committed 2nd echelon
     forces
C3.1: Drive YOC into desired chokepoints and restricted LOCs
C3.2: Disrupt movement of YOC forward corps and committed 2nd echelon
     forces

C3.3: Attack critical naval targets. Given that this supports CinC objective C3 (Deny use of national waterways for transport of military equipment), this has been rephrased to:

Attack naval transport ships which are capable of transporting military equipment on national waterways.

C4.1: Disrupt command and control of YOC Air/Land/Sea forces.

C4.2: Disrupt and Degrade YOC intelligence capability

C8.1: Disrupt command and control of YOC Air/Land/Sea forces

C8.2: Disrupt and Degrade YOC intelligence capability

C9.1: Deploy (Blue) forces capable of deterring YOC aggression

C9.2: Deploy (Blue) forces capable of supporting NEO

B.4 Fielded Forces

E4: Disrupt YOC Ground Scheme of Maneuver. Rephrased to: Disrupt Ground

     Scheme of Maneuver of YOC mechanized land forces

E7.1: Destroy strategic targets (pol-mil C2 and NBC/TBM capability)

E7.2: Disrupt YOC Weapons of Mass Destruction capability
E7.3: Destroy YOC Weapons of Mass Destruction capability

C Task Objectives

These had not been sorted by Checkmate at the time of writing. Categorization has been performed by the author of this document. Where a task objective falls into one category but services an air objective in another category, it is categorized under the former. For example, Destroy known NBC logistics support is categorized under National Infrastructure, even though it supports the air objective Disrupt YOC Weapons of Mass Destruction Capability, which is categorized under Key Production.

C.1 Leadership

C.1.1 YOC forces

Destroy key national headquarters
   Disrupt YOC reconnaissance/surveillance capabilities

Disable known NBC C3 capability

Disrupt C3I to frontline offensive units

Disrupt enemy control of air operations and (control of) air defenses
Destroy key command centers
Destroy air defense C3I
Disrupt key command and control nodes

Disrupt TBM C2

Neutralize airborne C3 capability

Disrupt key C3 nodes

Destroy key YOC LADS C2 facilities

Destroy command centers & comm nodes to disrupt YOC political and military direction of armed forces

Destroy army, corps, and division HQs or isolate them from subordi-

nate units

Sever LOCs in support of GCC ground scheme of maneuver. Rephrased to: Sever local command centers in support of GCC ground scheme of maneuver. Rephrased to:

Destroy coastal surface surveillance radar C2

Destroy key YOC LADS C2 facilities
Disrupt naval C3I
Disrupt Political Leadership of YOC Forces
Destroy Command and Control Facilities in the forward area

Destroy key YOC LOC's supporting forward troops Render ineffective C2 from YOC leadership to ballistic missiles

C.1.2 Blue forces

Provide battle management/C3I capability
Provide surveillance and reconnaissance of region
Destroy YOC intelligence collection/dissemination capability

Exploit CUWTF intel and targeting when tasked by the B/G C3

C.2 Key Production

Destroy known NBC research facilities
Destroy known NBC production facilities
Attrit military production capability

Disrupt electrical power supply to LADS and pol-mil C3I Disrupt/destroy military vehicles, armor, and artillery assembly & repair

facilities

Disrupt/destroy remaining POL production and storage

Disrupt/destroy remaining munitions and propellant production and
storage

Disrupt NBC production and storage capability
Disrupt nuclear reactor & fuel/ore processing
Destroy YOC mining capability
Disrupt YOC Electrical Power Distribution Destroy Naval Production
Capability

C.3 National Infrastructure

Disrupt YOC naval port operations
Destroy logistics/maintenance/supply areas
Disrupt SATCOM capability
Destroy known NBC logistics support
Destroy known NBC stockpiles and storage facilities
Disrupt electrical power grid
Destroy key naval communications nodes and EW/GCI
Destroy railroad transportation terminals and facilities
Destroy key bridges & tunnels

Disrupt operations along LOCS/chokepoints

Destroy POL storage/distribution facilities

C.4 National Population

None given.

C.5 Fielded Forces

C.5.1 YOC forces

Attrit follow on forces Attrit front line forces

Attrit enemy fighter/bomber force
Destroy key SAM sites

Destroy SAM sites threatening air operations throughout YOC Disrupt maneuver and momentum of committed forces with CAS

Neutralize YOC 240/170mm artillery

Destroy remaining mech & armor assets of 2nd echelon forces and strategic reserves

Disrupt YOC SAM/AAA sites along coastal corridors

Destroy Silkworm missile/sites and storage

Disrupt/destroy SAM/AAA/GCI affecting insertion corridors and drop zones

Render ineffective YOC target acquisition systems
Destroy landmines

Destroy key EW/GCI sites supporting YOC defense assets

Destroy coastal surface surveillance radar sites and C2. Divided into two: Destroy coastal surface surveillance radar sites is categorized here; destroying the radar’s C2 is categorized under Leadership

Neutralize YOC capability to utilize its aircraft and airfields. This is a higher level objective, since “capability to utilize” might encompass fuel, pilots, repair facilities, etc.

Destroy major YOC surface and sub-surface combatants and SOF insertion
platforms
Disrupt YOC Ground Forces With CAS
Destroy Tactical SAMs in YOC
Destroy Strategic SAMs in YOC

Destroy EW/GCI radars within 100 Km of the border

Disrupt YOC Fighter/Bomber Operations
Destroy Coastal Surveillance/Fire Control Radars
Destroy YOC SAM/AAA sites
Destroy Silkworm missiles/sites and storage

Destroy key EW/GCI sites supporting YOC defense assets

Disrupt YOC fighter capability

Disrupt/destroy SAMs on border, AAA near drop zones, and EW/GCI threatening CUWTF insertions

Disrupt YOC SAM/AAA sites along coastal corridors
Disrupt YOC bomber capability

Disrupt maneuver and momentum of committed forces with CAS Disrupt movement & maneuver into M-K corridor. Rephrased to: Disrupt

movement & maneuver of forces into M-K corridor Attrit forces in Wonchor corridor Disrupt/destroy 240mm/170mm artillery beyond the FSCL

Disrupt 240mm/170mm artillery inside the FSCL
Disrupt mechanized and armored formations between FLOT and FSCL
Destroy YOC SOF infiltration aircraft
Destroy major YOC surface and sub-surface combatants and SOF insertion
platforms

Destroy SAMs beyond 100km of the DMZ Assist in holding defense north of capital city by destroying YOC non-hardened

artillery and vehicles

Attrit Naval Bases and their Employment Capability Destroy enemy OCA assets

C.5.2 YOC supply, storage & repair facilities

Destroy TBM launchers, storage, and support
Disrupt TBM shipment/transshipment

Disrupt resupply of naval POL, munitions, & spares

   Destroy ammo storage
   Destroy naval ship repair & resupply facilities

Disrupt naval surface craft bases, support facilities, and naval C3I. This objective has been divided into 3 objectives: Disrupt naval surface craft bases and Disrupt support facilities are categorized under Fielded forces, while disrupting naval C3I is categorized under Leadership.

   Disrupt marshalling areas and sustainment capability beyond the FSCL
   Disrupt marshalling areas, supply, and corps C2 beyond FSCL

C.5.3 Blue forces

Employ passive defensive measures. Rephrased to: Assign and Provide military equipment which is capable of passive defense.

Provide ground/naval point defense for protection of Blue COGs from air & TBM attacks

Conduct CAS in support of the Rear Area Security Forces

Support B/G ground scheme of maneuver and deception plan through AI. Rephrased to: Provide AI to support B/G ground scheme of maneuver and deception plan.

C.6 High level objectives

The following were listed as task objectives, but are more appropriately categorized as either CinC objectives or Air Objectives.

Disrupt YOC ability to employ weapons of mass destruction (WMD). This is almost identical to Air Objective B8.2, and it belongs at the Air Objectives level.

Disrupt Employment of Weapons of Mass Destruction by YOC Forces Disrupt military activity in the area of the planned NEO Destroy command centers and comm nodes to disrupt YOC political di-

rection of armed forces

C.7 Unclassified

Destroy RADRELs – unknown abbreviation Employ SIGINT collection assets – unknown abbreviation Employ HVAA assets/Provide adequate HVAA protection – unknown ab-

breviation

D Goals

D.1 Goals identified

The goals identified are as follows:

• CinC Objectives

– Leadership ∗ Isolate;

∗ Neutralize;

∗ Protect

– Key Production ∗ Deny use;

∗ Neutralize

– National Infrastructure ∗ Deny use;

∗ Interrupt;

∗ Obtain;

∗ Permit use;

∗ Sustain

– National Population ∗ Deny use;

∗ Conduct NEO;

∗ Neutralize;

∗ Use

– Fielded forces ∗ Deter;

∗ Defend;

∗ Disrupt;

∗ Stop

• Air Objectives

– Leadership ∗ Isolate;

∗ Disrupt;

∗ Degrade;

∗ Destroy

– Key production:

∗ Disrupt;

∗ Destroy

– National Infrastructure ∗ Drive (i.e. compel to move);

∗ Disrupt;

∗ Attack;

∗ Deploy.

– National Population ∗ Disrupt^{4}

– Fielded forces ∗ Disrupt;

∗ Destroy

• Task Objectives

– Leadership ∗ Sever;

∗ Disrupt;

∗ Disable;

∗ Neutralize;

∗ Render ineffective;

∗ Destroy;

∗ Exploit;

∗ Provide.

– Key production:

∗ Disrupt;

∗ Attrit;

∗ Destroy

– National Infrastructure ∗ Disrupt;

∗ Destroy

– National Population No task objectives specified.

– Fielded forces ∗ Disrupt;

∗ Attrit;

∗ Neutralize;

∗ Destroy;

∗ Provide;

∗ Conduct CAS.

4No air objectives are specified for national population. This goal is specified because neutralize appears in the CinC objectives, and so an analogy can be drawn with the Leadership and Key Production objectives.

D.2 Ordering and definitions of goals

Table 4 lists identified goals in a suggested order, from the least effect to the greatest. Definitions of each goals are also suggested in order to support the ordering.

D.3 Goals and suggested opposite goals

Table 5 lists goals, and suggests opposite goals. Note that these opposite goals are intended to be the positive antithesis of each destructive goal, as opposed to simply maintaining the status quo by preventing the desctructive goals from being achieved.

Goal  Definition

Leadership

Disrupt  Temporarily reduce effectiveness
Degrade  Permanently reduce effectiveness
Isolate/Sever  Reduce effectiveness to zero until further notice

Neutralize/Disable/ Render ineffective

Reduce effectiveness to zero until further notice

Destroy  Permanently reduce effectiveness to zero

Key production

Disrupt  Temporarily reduce production
Attrit  Permanently reduce production
Destroy  Permanently stop production

Infrastructure

Disrupt  Reduce the flow
Destroy  Permanently stop the flow

National Population

Disrupt  Temporarily reduce effectiveness

Fielded forces: location

Deter Make it dangerous (politically or militarily) for YOC forces to advance or attack

Defend  Reduce the effectiveness of YOC forces when attacking
Drive  Compel YOC forces to travel on certain routes
Stop  Prevent YOC forces from advancing
Expel  Compel YOC forces to retreat

Fielded forces: capability

Disrupt Temporarily reduce military capability of YOC force

Attrit Permanently reduce military capability of YOC force

Neutralize Temporarily reduce YOC force’s military capability to zero

Destroy Permanently reduce YOC force’s military capability to zero

Table 4: Goals, ordered by effect

Goal  Opposite

Leadership

Disrupt  Put into order
Degrade  Improve
Isolate  Connect
Neutralize  Restart
Destroy  Create

Key production

Disrupt  Put into order
Destroy  Create

Infrastructure

Disrupt  Put into order
Interrupt  Restart
Deny use  Permit use
Destroy  Create

National Population

Disrupt  Put into order

Fielded forces: location

Deter  Encourage
Defend  Attack
Drive  Open new routes
Stop  Advance
Expel  Hold position

Fielded forces: capability

Disrupt  Put into order
Attrit  Build up
Neutralize  Restart
Destroy  Create

Table 5: Goals and their opposites

E Centers of Gravity

26 Figure 2: CoGS: Leadership

27 Figure 3: CoGS: Key Production 1: General

28 Figure 4: CoGS: Key Production 2: Military

29 Figure 5: CoGS: National Infrastructure 1: Transportation

30 Figure 6: CoGS: National Infrastructure 2: Communications

31 Figure 7: CoGS: National Infrastructure 3: Fuel & Power

32 Figure 8: CoGS: National Population

33 Figure 9: CoGS: Fielded Forces 1: Special Weapons

34 Figure 10: CoGS: Fielded Forces 2: Land forces

35 Figure 11: CoGS: Fielded Forces 3: Naval forces

36 Figure 12: CoGS: Fielded Forces 4: Air forces

37 Figure 13: CoGS: Fielded Forces 5: Logistics

38

F Constraints

39

Geographical

US citizens  in the region
US forces  in the region

 YOC aggression  against B/G

YOC forces  north of Desired Area

       B/G

(a geographical constraint is implicit in the CoG)

Temporal

YOC economic centers

to slow the flow of military resources i.e. until the flow of military resources is sufficiently slow

Categorical

YOC political control
YOC military control

YOC armed forces

YOC political control
YOC military control
primary electrical sources
petroleum products
national television systems
national radio systems
national computer systems

YOC population

national roadways
national railways
national waterways

Transport of military equipment

national roadways
national railways
national waterways

Transport of goods

national communications networks
national television systems
national radio systems
national computer systems

Transmit military information

YOC petroleum production  supporting warfighting effort

chemical products

for the manufacture of weapons of mass destruction.

national waterways  for Blue shipping.
national transportation networks  leading to YOC economic centers.

Blue forces  Against Weapons of Mass Destruction

Geographical Temporal Categorical

YOC political control
YOC military control
electrical facilities
petroleum production
chemical production?

YOC armed forces

YOC political control
YOC military control
electrical production
petroleum production
 chemical production

YOC population

naval transport ships

which are capable of transporting military equipment on national waterways.

command and control  YOC Air/Land/Sea forces

Capability

(Blue) forces  capable of deterring YOC aggression
(Blue) forces  capable of supporting NEO

Table 6: Constraints: Air Objectives Level